# One missing piece in Vision and Language: A Survey on Comics Understanding

### v1

@@ 8.2 Modification
In the realm of comics, modification tasks such as image inpainting and image editing play a crucial role in streamlining the storytelling process and enhancing creative expression.
8.2.1 Image Inpainting and Editing
@@
One innovative tool in this space is CodeToon, presented by [169]. CodeToon is a comic authoring tool that facilitates a code-driven storytelling process. It leverages two mechanisms: story ideation from code through metaphorical interpretation and automatic comic generation from these ideated stories. Significant attention has been given to the interface design to address common challenges faced by comic creators, such as generating stories directly from code snippets and seamlessly transitioning these stories into comic format.
Another notable contribution is ComicScript by [170], also known as interactive data comics [^11] Accessible at https://interactivedatacomics.github.io/.. ComicScript is a suite of operations designed in collaboration with practitioners to create and modify comics. These operations allow for the addition and removal of panels, support branching narratives, perspective changes, and detailed exploration. They also facilitate data manipulation and interaction within the comics, enhancing personalization and reader engagement. To validate the utility of ComicScript, professional illustrators, designers, and data comics enthusiasts were recruited to craft interactive comics. Their experiences provided valuable insights into the authoring workflow and the potential of interactive comics as a storytelling medium.
8.3 Future Tasks
8.3.1 Personalized Image retrieval
Definition: Personalized Image Retrieval (PIR) focuses on developing person-centric models that can efficiently retrieve specific comic panels based on a compound query involving both image and text inputs [171]. PIR aims to identify and retrieve images or videos that correspond to a compound query, which typically includes an image of a person’s face combined with a textual scene or action description.
@@ 8.3.1 Personalized Image retrieval
9 Layer 4: Advanced Visual-Language Understanding
9.1 Understanding
This layer investigates tasks that require a deeper integration of visual elements and language, advancing beyond basic identification or retrieval to more complex forms of comprehension and reasoning within comics.
9.1.1 Visual Question Answering
@@
The first notable attempt in the domain of the comic is by [172] with ComicVQA. However, their approach, more akin to retrieval based on specific “Doraemon” comics datasets, diverges from the typical VQA framework. Later, Sumi et al. [173] presented a system, ComicsQA, with a unique twist, converting user QA interactions into comic stories, reflecting the user’s situation and offering solutions.
The creation of a comprehensive VQA dataset specifically for comics, encompassing single-panel, single-page, and multi-page formats, remains an open challenge. In a recent survey [174], Zeng et al. show the massive work that has been carried out on VQA on multi-modal data across different mediums (movies, comics, etc.) and Multimodal Machine Comprehension [175, 176]. This, together with multiple datasets across various mediums [177, 178, 179, 180, 181, 181] are laying the groundwork for advancing VQA also in comics.
9.2 Future Tasks
@@
9.2.1 Visual Entailment
Definition: Visual entailment (VE) in comics is an area yet to be deeply explored. This task involves assessing whether a given image semantically entails the accompanying text. The challenge lies in interpreting the visual narrative in conjunction with textual elements to ascertain semantic consistency or correlation. It can be seen as a simplified version of the text-closure task where we aim at correctly classifying whether the image is represented by the given text, however, depending on the provided text, VE could quickly lead to complex scenarios.
9.2.2 Visual Dialog
@@ 9.2.3 Visual Reasoning
10 Layer 5: Generation and Synthesis
@@
10.1 Generation
The field of Generation in comics has seen lots of attraction on Manga style comics rather than American ones. Their popularity has attracted the research community towards creating comics from various media such as video, images, and text descriptions.
10.1.1 Comics generation from other media.
@@ 10.1 Generation
One of the earliest attempts at comic generation from other media is the “Comic Live Chat” by Matsuda et al. [182], which transformed video meetings into comics. This approach involved selecting keyframes and rendering the dialogues of meeting participants in a comic format. Tanapichet et al. [183] proposed a system for creating comic strips from cartoon animations using optical flow techniques, demonstrating a novel approach to narrative translation from one visual medium to another. Hoashi et al. [184] worked on creating manga previews by detecting panels and text balloons, while in [185, 186] authors explored generating manga faces and caricatures from real images. Cao et al. [187] and Herranz et al. [188] introduced methods for comic-like video summarization. These systems developed pipelines for layout proposal, image reconfiguration, and balloon creation, treating comic creation as a series of steps from keyframe selection to final panel layout. On the same task, to automatically create comics from video content, Jing et al. [189] pioneered video transformation using TV show subtitles and a speaker detection algorithm to generate comic layouts optimized through an energy-based metric. This approach was a significant step forward compared to earlier methods like [190].
@@
Giving it more time, a groundbreaking approach by Pesko et al. [191] employed neural style transfer, creating visually compelling comic representations from video-sourced data [^12] GitHub repository: https://github.com/maciej3031/comixify. While their startup has since become closed-sourced, their initial approach signifies a significant leap in comic generation technology. Yang et al. [192] presented a system to generate comics from movies, employing a multi-page layout framework and emotion-aware balloon generation, showcasing the potential of comprehensive systems in comic generation. To extract keyframes from movies, they used GIST [193] similarity between two frames. To detect the speaker, they perform lip motion analysis and use that to associate balloons with talking characters. Finally, the movie frames are rendered in a comic-layout page after the adaption to comic style. Moreover, in [194] authors represent conversational videos to comics “fixed-layout” representation, developing a system that learns guidance field which provides a prior prediction of the possible positions of word balloons while making the word balloons not overlap with other nonverbal information (e.g. hand gestures, visual clues in the background, etc). Their input is raw video, thus the text dialog is captured from audio and rendered in the location chosen from their method.
Comics have been used also in education [195, 196, 197] to teach and explain in an easier way different concepts. An example of this is [198] where authors proposed a system that crafts data stories from a collection of user-created charts, using a comic-style panel to imply the underlying sequence and logic of data-driven narratives.
10.1.2 Image-to-Text generation
@@
Motivated by supporting the Blind or Low Vision community (BLV) - also known as People with Visual Impairements (PVI) - Ramaprasad et al. [199] developed a two-step method to create natural language descriptions of comic strips. This method first extracts information about panels, characters, and text using computer vision techniques, followed by the use of a multimodal large language model (MLLM) to generate descriptions. This work stands as one of the few promoting MLLM specifically for comic strip captioning, despite using “out of the box” architectures like Grounding DINO [40] to extract arbitrary elements, CLIP [14] to match character images with name-descriptions, and LLaVa [200] as MLLM.
@@
Vivoli et al. [72] proposed an advanced version of the text-cloze task with image-to-text generation, increasing the challenge level. They adopted the VL-T5 model [16] using two configurations with similar performances but different model sizes: with BLIP-2 visual backbone (1.3B) or a ResNet-50 fine-tuned with SimCLR on comics (0.2B). They surpassed previous state of the art in the standard “text-cloze” task, and propose benchmarks in the generation version of the task. Authors in [77] presented a similar task but incorporated character descriptions and transcriptions, emphasizing a language model-centric approach. Guo et al. [201] introduced the multi-modal manga complement task. This innovative task combines visual sequences from comic pages with corresponding text dialogues, challenging the model to complete the narrative appropriately. They propose an effective method with CLIP as Feature Encoding, a custom cross-attention module called Fine-grained Visual Prompt Generation, and a transformer-based encoder-decoder architecture called Dialog complement. In a recent work [41], authors proposed a Relationformer-based architecture that generates dialog for manga pages. Their approach includes detection, matching, and sorting algorithms to produce an ordered dialog that aligns with the visual narrative. However, all these works are limited by the dialog appearing in the comic. No work has explored producing a multi-page description of the scene, the actions, and the surrounding environment of every single panel.
10.1.3 Text-to-Image generation
@@
In the field of Text-to-Image generation, one of the earliest endeavors was by Jin et al. [202], who successfully created anime-style character faces using a GAN trained on a curated dataset of anime faces, showcasing the potential of GANs in generating stylized images. Inoue et al. [68] introduced novel methods for domain transfer and pseudo-labeling in image generation, marking significant advancements in the area. StoryGAN [203] emerged as a groundbreaking initiative in story visualization, employing a sequential conditional GAN framework equipped with a context encoder to dynamically track story flow and a story-level discriminator. Enhancements to StoryGAN have included a dual learning framework to bolster semantic consistency between stories and images by refining captions [204] and a recursive architecture to handle structured text inputs via constituency parsing trees [205]. In [206], a deep learning pipeline was proposed for generating synthetic graphic novels using GPT-2 for text and StyleGAN2 for image synthesis, trained on a large collection of manga pages. Proven-Bessel et al. [207] developed a text-to-image GAN for generating single-panel comics in the style of Dilbert, showcasing new possibilities for comic generation from text. Another approach involved FastGAN [208], enhanced with a condition vector for generating high-quality images from small datasets, such as manga faces drawn from Osamu Tezuka’s works [209], which showed improved FID scores compared to the original FastGAN. In [210], a multilingual text-to-image model was developed for creating webtoons, demonstrating the adaptability of GANs to multilingual contexts. However, despite various training experiments, the webtoon fine-tuned GAN underperformed compared to fine-tuning Korean MSCOCO, with the latter still producing abstract images that noticeably differed from authentic comics.
Significant progress has also been observed in story visualization through diffusion models. Innovations like StoryDALL-E [211] employ autoregressive Transformers for extensive model tuning and adaptive adjustments. Everaert et al. [212] discovered the ability to adapt Stable Diffusion to various styles, including comics, by adjusting the initial latent tensor, showing how diffusion models can contribute to comic generation. Despite diffusion models, both on pixel and on latent space, being not good at generating text, the comic panels generated are in line with COMIC data distribution. The topic of story visualization as a subfield of text-to-image generation is very active, moving also towards integration of more powerful LLMs [213], with the main challenges being the consistency of characters and style across multiple images (panels).
Lastly, as we saw previously for cross-modal retrieval tasks, using existing models as building blocks to automatically annotate new data is a good option, especially when the task doesn’t exist or annotations are expensive. In [214], authors used an LLM (ChatGPT) to generate storylines and dialogue of a “possible continuation” of the well-known manga One Piece. Then, they fine-tuned stable diffusion with LoRA and ControlNet to adapt to the manga layout, colorization, and character aesthetic. They finally generated the comic using the fine-tuned stable diffusion, rendered it in panels, and added the LLM dialogs to balloons and narration boxes.
10.1.4 Sound Generation
Definition: Sound generation in the comics domain stands for generating sound effects for a single panel, or in the transaction from one panel to another.
In [215], authors proposed a system that automatically narrates the input comic page of “Clubhouse Rascals” by synthesizing realistic speeches for each character. The synthesized speeches carry the identity properties (gender and age) and emotional conditions (e.g., happy, sad, angry) of the characters inferred from the comic input. They perform a series of subsequential steps including detection of panels, speech balloons, text, character, and character’s face, as well as character re-identification, text-character association, and dialog generation. They classify every character’s face with gender, age, and emotional state, and generate the comic speech synthesis given all this information as input to a waveform reconstruction. They used the Manga109 dataset for all experiments.
In a later study [216], authors explored whether having scene description and sound effects could improve the user experience of comics for People with Visual Impairments (PVI), discovering that both types are welcomed with the first (scene description) helping fr concentration and understanding of the situation, and the second one (sound effect) perceived as more immersive and realistic book-reading experience.
10.2 Synthesis
@@
The group of Synthesis tasks, compared to the “generation” one, refers to the creation of more complex, structured, and often temporally extended outputs, such as creating a full graphic novel from a complex storyline or merging multiple elements (characters, scenes, dialogues) in a consistent way across multiple pages.
10.2.1 3D generation from images
@@
While there are advancements in using 3D character assets for tasks like pose estimation [217], re-targetting [218, 219], and character reposing [220], the field lacks scalable training resources due to the unavailability of suitable 3D character assets. In light of these issues, Chen et al. [221] addressed this gap by formalizing the stylized reconstruction task with the introduction of the AnimeRecon benchmark and the Vroid dataset. The Vroid dataset, in particular, provides a wealth of 3D assets for scalable training in this domain. A significant part of their work involves solving the challenge of contour removal from illustrations, which is a crucial step in achieving accurate 3D reconstruction from stylized comic images.
While there are numerous works on 3D reconstruction from line drawings in general [222, 223], the specific application in comics is underexplored, likely due to the absence of suitable datasets and ground truth references for comics.
It is worth noting that the exploration of 3D reconstruction from comic images holds great potential. It not only enriches the visual experience of comics but also opens up possibilities for animations, interactive media, and virtual reality applications within the comic domain.
10.2.2 Video generation

### v2

@@ 8.2. Modification
In the realm of comics, modification tasks such as image inpainting and image editing play a crucial role in streamlining the storytelling process and enhancing creative expression (Fig. 12).
8.2.1. Image Inpainting and Editing
@@
In the realm of comic inpainting and editing, CodeToon (Suh and Zhao, 2022) offers a code-driven storytelling tool that automatically generates comics from code snippets through metaphorical interpretation. It simplifies the transition from story ideation to comic format, addressing common challenges faced by comic creators. ComicScript (Wang et al., 2021) provides a suite of operations for interactive comic creation, allowing users to add/remove panels, change perspectives, and support branching narratives. It also integrates data manipulation to enhance engagement. Professional creators validated its potential by crafting interactive comics, showcasing the tool’s ability to personalize storytelling. Both tools highlight the growing role of computational methods in enhancing comic creation and editing workflows.
8.3. Future Tasks
8.3.1. Personalized Image retrieval
Definition: Personalized Image Retrieval (PIR) focuses on developing person-centric models that can efficiently retrieve specific comic panels based on a compound query involving both image and text inputs (Korbar and Zisserman, 2022). PIR aims to identify and retrieve images or videos that correspond to a compound query, which typically includes an image of a person’s face combined with a textual scene or action description.
@@ 8.3.1. Personalized Image retrieval
9. Layer 4: Advanced Visual-Language Understanding
In the fourth layer, we propose in the Layer of Comics Understanding a set of tasks related to multimodal understanding, which comprehends often reasoning through a text-rich image given a text prompt and one or more images as input. However, the only task that has been explored in comics already is Visual Question Answering, which under its umbrella groups different definitions of VQA in comics.
9.1. Understanding
Figure 13. Illustration of Understanding tasks.
This layer investigates tasks that require a deeper integration of visual elements and language, advancing beyond basic identification or retrieval to more complex forms of comprehension and reasoning within comics (Fig. 13).
9.1.1. Visual Question Answering
@@
The first notable attempt in the domain of the comic is by (Moriyama et al., 2016) with ComicVQA. However, their approach, more akin to retrieval based on specific “Doraemon” comics datasets, diverges from the typical VQA framework. Later, Sumi et al. (Sumi, 2017) presented a system, ComicsQA, with a unique twist, converting user QA interactions into comic stories, reflecting the user’s situation and offering solutions. The creation of a comprehensive VQA dataset specifically for comics, encompassing single-panel, single-page, and multi-page formats, remains an open challenge. In a recent survey (Zeng et al., 2020), Zeng et al. show the massive work that has been carried out on VQA on multi-modal data across different mediums (movies, comics, etc.) and Multimodal Machine Comprehension (Sahu et al., 2021b, a). This, together with multiple datasets across various mediums (Singh et al., 2019; Biten et al., 2019; Mathew et al., 2022; Vivoli et al., 2022; Tanaka et al., 2023, 2023) are laying the groundwork for advancing VQA also in comics.
9.1.2. Visual Entailment
Definition: Visual entailment (VE) in comics is an area yet to be deeply explored. This task involves assessing whether a given image semantically entails the accompanying text. The challenge lies in interpreting the visual narrative in conjunction with textual elements to ascertain semantic consistency or correlation. It can be seen as a simplified version of the text-closure task where we aim at correctly classifying whether the image is represented by the given text, however, depending on the provided text, VE could quickly lead to complex scenarios. Recently, (Hu et al., 2024) investigated the performances of Multimodal models on comics with contradictory narratives, where each comic consists of two panels that create a humorous contradiction. The paper introduces the YesBut benchmark and assesses Multimodal Models’ limited capabilities in this reasoning task.
9.2. Future Tasks
@@
9.2.1. Visual Dialog
@@ 9.2.2. Visual Reasoning
10. Layer 5: Generation and Synthesis
@@
10.1. Generation
Figure 14. Illustration of Generation tasks.
The field of Generation in comics has seen lots of attraction on Manga style comics rather than American ones. Their popularity has attracted the research community towards creating comics from various media such as video, images, and text descriptions (Fig. 14).
10.1.1. Comics generation from other media.
@@ 10.1. Generation
One of the earliest attempts at comic generation from other media is the “Comic Live Chat” by Matsuda et al. (Matsuda et al., 2010), which transformed video meetings into comics. This approach involved selecting keyframes and rendering the dialogues of meeting participants in a comic format. Tanapichet et al. (Tanapichet et al., 2011) proposed a system for creating comic strips from cartoon animations using optical flow techniques, demonstrating a novel approach to narrative translation from one visual medium to another. Hoashi et al. (Hoashi et al., 2011) worked on creating manga previews by detecting panels and text balloons, while in (Deng et al., 2010; Chang and Cheng, 2011) authors explored generating manga faces and caricatures from real images. Cao et al. (Cao et al., 2012) and Herranz et al. (Herranz et al., 2012) introduced methods for comic-like video summarization. These systems developed pipelines for layout proposal, image reconfiguration, and balloon creation, treating comic creation as a series of steps from keyframe selection to final panel layout. On the same task, to automatically create comics from video content, Jing et al. (Jing et al., 2015) pioneered video transformation using TV show subtitles and a speaker detection algorithm to generate comic layouts optimized through an energy-based metric. This approach was a significant step forward compared to earlier methods like (Wang et al., 2012). Giving it more time, a groundbreaking approach by Pesko et al. (Pesko et al., 2018) employed neural style transfer, creating visually compelling comic representations from video-sourced data [^9] GitHub repository: https://github.com/maciej3031/comixify. While their startup has since become closed-sourced, their initial approach signifies a significant leap in comic generation technology. Yang et al. (Yang et al., 2021) presented a system to generate comics from movies, employing a multi-page layout framework and emotion-aware balloon generation, showcasing the potential of comprehensive systems in comic generation. To extract keyframes from movies, they used GIST (Oliva and Torralba, 2001) similarity between two frames. To detect the speaker, they perform lip motion analysis and use that to associate balloons with talking characters. Finally, the movie frames are rendered in a comic-layout page after the adaption to comic style. Moreover, in (Zhang et al., 2023) authors represent conversational videos to comics “fixed-layout” representation, developing a system that learns guidance field which provides a prior prediction of the possible positions of word balloons while making the word balloons not overlap with other nonverbal information (e.g. hand gestures, visual clues in the background, etc). Their input is raw video, thus the text dialog is captured from audio and rendered in the location chosen from their method. Comics have been used also in education (Suh et al., 2021; Lima et al., 2022; Castro et al., 2022) to teach and explain in an easier way different concepts. An example of this is (Zhao et al., 2021) where authors proposed a system that crafts data stories from a collection of user-created charts, using a comic-style panel to imply the underlying sequence and logic of data-driven narratives.
10.1.2. Image-to-Text generation
@@
Motivated by supporting the Blind or Low Vision community (BLV) - also known as People with Visual Impairements (PVI) - Ramaprasad et al. (Ramaprasad, 2023) developed a two-step method to create natural language descriptions of comic strips. This method first extracts information about panels, characters, and text using computer vision techniques, followed by the use of a multimodal large language model (MLLM) to generate descriptions. This work stands as one of the few promoting MLLM specifically for comic strip captioning, despite using “out of the box” architectures like Grounding DINO (Liu et al., 2023c) to extract arbitrary elements, CLIP (Radford et al., 2021) to match character images with name-descriptions, and LLaVa (Liu et al., 2023b) as MLLM. Vivoli et al. (Vivoli et al., 2024d) proposed an advanced version of the text-cloze task with image-to-text generation, increasing the challenge level. They adopted the VL-T5 model (Cho et al., 2021) using two configurations with similar performances but different model sizes: with BLIP-2 visual backbone (1.3B) or a ResNet-50 fine-tuned with SimCLR on comics (0.2B). They surpassed previous state of the art in the standard “text-cloze” task, and propose benchmarks in the generation version of the task. Authors in (Agrawal et al., 2023) presented a similar task but incorporated character descriptions and transcriptions, emphasizing a language model-centric approach. Guo et al. (Guo et al., 2023) introduced the multi-modal manga complement task. This innovative task combines visual sequences from comic pages with corresponding text dialogues, challenging the model to complete the narrative appropriately. They propose an effective method with CLIP as Feature Encoding, a custom cross-attention module called Fine-grained Visual Prompt Generation, and a transformer-based encoder-decoder architecture called Dialog complement. In a recent work (Sachdeva and Zisserman, 2024), authors proposed a Relationformer-based architecture that generates dialog for manga pages. Their approach includes detection, matching, and sorting algorithms to produce an ordered dialog that aligns with the visual narrative. However, all these works are limited by the dialog appearing in the comic, and in the context of describing a comic, not only the dialog is important, but also the scene description. Recent works, to fill this gap, have approached the task by exploring single panel captioning. In (Rigaud et al., 2024), the authors explored prompt engineering with contextual information for closed-sourced Multimodal LLMs aiming at generating an accurate text description of a single panel and subsequently of the full story. They successfully demonstrated to retain many of the important details in the caption, but the exploration was limited to a couple of story panels. In (Vivoli et al., 2024b), authors took a step further proposing a new metric for comic panel captioning called “attribute retaining metric” (ARM) which assesses whether all the objects and attributes of the panel have been identified. They propose a pipeline with open-source Multimodal LLMs and zero-shot grounding models for generating dense and grounded captions for more than 1M panels. However, in the ARM metric only the attributes are taken into consideration, which identifies the need to consider both locations and object/attribute names. Finally, (Hu et al., 2024) focuses on comics with contradictory narratives, where each comic consists of two panels that create a humorous contradiction, and introduces the “YesBut” benchmark, which comprises tasks of varying difficulty aimed at assessing MLLMs capabilities in recognizing and interpreting these comics.
10.1.3. Text-to-Image generation
@@
In the field of Text-to-Image generation, early work by Jin et al. (Jin et al., 2017) utilized GANs to generate anime-style character faces, demonstrating the potential of GANs for stylized image generation. Inoue et al. (Inoue et al., 2018) advanced the field with domain transfer and pseudo-labeling techniques for image generation. StoryGAN (Li et al., 2018a) was a groundbreaking development in story visualization, employing a sequential conditional GAN with a context encoder to track story flow and a story-level discriminator. Later improvements included DuCo-StoryGAN (Maharana et al., 2021), which used a dual learning framework for better semantic consistency, and VLStoryGAN (Maharana and Bansal, 2021), which incorporated recursive architectures to handle structured text inputs. Melistas et al. (Melistas et al., 2021) proposed a pipeline for generating synthetic graphic novels, using GPT-2 for text and StyleGAN2 for image synthesis, trained on a large manga collection. Proven-Bessel et al. (Proven-Bessel et al., 2021) developed ComicGAN, a text-to-image GAN for generating single-panel comics in the style of Dilbert, while FastGAN (Liu et al., 2021b), enhanced for small datasets, was applied to manga faces (Hiruta et al., 2022) with improved FID scores. Yu et al. (Yu et al., 2023) introduced a multilingual text-to-image model for webtoon generation, showing the flexibility of GANs across languages. However, even fine-tuning on Korean MSCOCO resulted in abstract images that differed from authentic comics. Diffusion models have also made strides in story visualization. StoryDALL-E (Maharana et al., 2022) used autoregressive Transformers for improved model tuning. Everaert et al. (Everaert et al., 2023) adapted Stable Diffusion for comic generation by tweaking the initial latent tensor, showing how diffusion models can be tailored for comic styles. Despite limitations in text generation, diffusion models successfully generated comic panels aligned with COMIC data distributions. The consistency of characters and style across multiple images, especially in storytelling contexts, remains a key challenge (He et al., 2024). Recent work by Jin et al. (Jin and Song, 2023) leveraged large language models (LLMs) and fine-tuned Stable Diffusion to generate new manga content. Using ChatGPT to generate storylines and dialogues for a hypothetical continuation of the manga One Piece, they fine-tuned Stable Diffusion with LoRA and ControlNet to match the layout, style, and aesthetic of the original manga. The final product included LLM-generated dialogue integrated into comic panels. The subfield of story visualization within text-to-image generation is rapidly evolving, driven by the integration of powerful LLMs and new image synthesis techniques. Key challenges include maintaining character and style consistency across multiple panels and developing unified benchmarks to track advancements across the field.
10.2. Synthesis
Figure 15. Illustration of Synthesis tasks.
@@ 10.2. Synthesis
The group of Synthesis tasks, compared to the “generation” one, refers to the creation of more complex, structured, and often temporally extended outputs, such as creating a full graphic novel from a complex storyline or merging multiple elements (characters, scenes, dialogues) in a consistent way across multiple pages (Fig. 15).
10.2.1. 3D generation from images
@@
While there are advancements in using 3D character assets for tasks like pose estimation (Kim et al., 2019), re-targetting (Khungurn, 2022; Kim et al., 2021), and character reposing (Lin et al., 2022), the field lacks scalable training resources due to the unavailability of suitable 3D character assets. In light of these issues, Chen et al. (Chen et al., 2023a) addressed this gap by formalizing the stylized reconstruction task with the introduction of the AnimeRecon benchmark and the Vroid dataset. The Vroid dataset, in particular, provides a wealth of 3D assets for scalable training in this domain. A significant part of their work involves solving the challenge of contour removal from illustrations, which is a crucial step in achieving accurate 3D reconstruction from stylized comic images. While there are numerous works on 3D reconstruction from line drawings in general (Lun et al., 2017; Sanghi et al., 2023), the specific application in comics is underexplored, likely due to the absence of suitable datasets and ground truth references for comics. It is worth noting that the exploration of 3D reconstruction from comic images holds great potential. It not only enriches the visual experience of comics but also opens up possibilities for animations, interactive media, and virtual reality applications within the comic domain.
10.2.2. Video generation